A stone of mass 0.15 kg is fired from a catapult. The velocity of the stone changes from 0 to 4.0 m/s in 0.50 s. What is the average resultant force acting on the stone while it is being fired?
1
Expert's answer
2020-12-08T10:46:32-0500

We can find the average resultant force acting on the stone while it is being fired from the impluse-momentum change equation:


"F\Delta t=m\Delta v,""F=\dfrac{m\Delta v}{\Delta t},""F=\dfrac{0.15\ kg\cdot(4.0\ \dfrac{m}{s}-0\ \dfrac{m}{s})}{0.50\ s}=1.2\ N."

Answer:

"F=1.2\ N."